# Generalized Distribution Prediction for Asset Returns

### v1

@@ Abstract
We present a novel approach for predicting the distribution of asset returns using a quantile-based method with Long Short-Term Memory (LSTM) networks. Our model is designed in two stages: the first focuses on predicting the quantiles of normalized asset returns using asset-specific features, while the second stage incorporates market data to adjust these predictions for broader economic conditions. This results in a generalized model that can be applied across various asset classes, including commodities, cryptocurrencies, as well as synthetic datasets. The predicted quantiles are then converted into full probability distributions through kernel density estimation, allowing for more precise return distribution predictions and inferencing. The LSTM model significantly outperforms a linear quantile regression baseline by 98% and a dense neural network model by over 50%, showcasing its ability to capture complex patterns in financial return distributions across both synthetic and real-world data. By using exclusively asset-class-neutral features, our model achieves robust, generalizable results.
@@
Machine Learning, Distribution Prediction, Asset Return Forecasting, Long Short-Term Memory (LSTM) Networks
@@ 1 Introduction
Every day, an almost uncountable number of transactions take place in capital markets, with each transaction generally aimed at generating a positive return on investment. If one could predict future financial asset prices accurately, the potential for gains would be nearly limitless. Even achieving a prediction accuracy slightly above 50% could potentially turn someone into a billionaire [35]. However, predicting returns is an extraordinarily difficult problem. Nobel laureate Eugene Fama famously argued that future returns are inherently random and, therefore, practically impossible to predict [11]. This difficulty is further supported by the fact that many funds consistently fail to outperform the S&P 500, a common benchmark for returns [13, 19].
@@
In finance, logarithmic asset returns are often assumed to follow a normal distribution, a famous example is the Black-Scholes model which is used to find the fair value of option derivatives [3]. However, in practice, this assumption frequently does not hold [26]. Returns often display non-normal characteristics, such as skewness or heavy tails. Consequently, models that rely on normality assumptions can lead to suboptimal or even incorrect results [30]. To address this issue, some approaches generalize assumptions or adopt alternative distributions, such as the Student’s T distribution, which accommodates skewness, kurtosis, and heavy tails [1]. More recently, a new approach has emerged, namely to model the distribution directly [2]. This approach offers several key advantages. A more accurate model of the return distribution could improve derivative pricing, inform risk management, and enhance portfolio optimization.
@@
In this paper, we propose a model that uses a distributional approach to predict asset returns. Our work extends the model presented in [2], which predicts the distribution of stock returns. Instead of predicting the return itself, our model predicts the quantiles of the return distribution. Although quantile regression is not a novel concept, we convert these quantiles into a full distribution using kernel density estimation. This enables a more precise estimation of the return distribution over a given time period. The model is trained on a dataset of asset returns and general market variables, ensuring that each feature is asset-class neutral, making the approach applicable across different asset classes.
This paper contributes to the literature by creating a more generalizable distribution prediction model, utilizing only asset-neutral features and allowing for the consideration of multiple return types. The results are compared to a baseline linear quantile regression model and the original approach in [2]. We find that our generalized approach outperforms the baseline by 121% on real-world data, while slightly underperforming by 1% on synthetic data. Compared to the original approach, our model outperforms on both real-world and synthetic data by 50% and 64%, respectively.
@@ 2 Literature Review
2.1 Asset Return Prediction
The field of asset return prediction using machine learning and artificial intelligence has expanded rapidly in recent years. The sheer volume of published studies makes it increasingly difficult to conduct a thorough review of the existing literature. As noted by Dessain, the number of academic studies has grown at an exponential rate, creating challenges for researchers attempting to synthesize existing knowledge [8]. He observes that conducting a comprehensive review of the literature on machine learning in finance is "becoming more and more challenging, if feasible at all," due to the rapid influx of studies [8]. Similarly, Strader et al. [31] highlighted the difficulty in comparing studies because of the vast diversity of approaches and benchmarks used across research.
In addition to the sheer volume of publications, the complexity and diversity of machine learning models used in asset return prediction further complicates synthesis. The review by Shah et al. underscores the variety of methods employed, including ARIMA [20, 9, 25], LSTM [28, 34, 21], CNN [5, 15, 17], and hybrid models [32, 24, 14]. This wide range of methodologies makes it challenging to compare results across studies [29]. This issue aligns with Dessain’s observation that conducting a comprehensive review is exceptionally difficult [8].
Another factor that hinders a thorough and consistent literature review is the variation in data used across studies [8]. Differences in asset classes, individual assets, time periods, and time intervals, often chosen arbitrarily, complicate meaningful comparisons between methodologies. To mitigate this, our paper uses publicly available data, ensuring that our results can be easily compared to other studies.
While several studies have attempted to address these challenges [27, 18, 4], they too struggle to keep pace with the rapid rate of publication and face the same issues described above. These efforts are important in their own right, but the complexities of the field continue to pose significant challenges for comprehensive reviews.
2.2 Distribution Prediction
Distribution prediction is a relatively novel approach in the machine learning literature. One of the few papers to focus on this is [2], where the authors propose a model that predicts quantiles of the return distribution rather than the return itself. This approach aligns with the quantile regression framework discussed by Liu [23]. The concept has been further advanced by models like the Conditional Quantile Variational Autoencoder (CQVAE), introduced by Yang et al. [33], which employs neural networks to model conditional quantiles. Together, these studies demonstrate the potential of distribution prediction and highlight its advantages over more traditional methods, which are often constrained by certain assumptions about return distributions.
Despite these advances, there remains significant room for improvement, and this is where our work contributes. While [2] predicts the 22-day ahead quantiles of returns for a specific asset, our model leverages Long Short-Term Memory (LSTM) networks to predict quantiles for an $n$-day period, where $n$ corresponds to both the input sequence length and the forecast window.
Our approach also differs in the type of input features used. The model in [2] incorporates 194 stock-specific characteristics along with macroeconomic variables to predict return quantiles. In contrast, our model employs asset-neutral features, drawing on the generalization techniques from the CQVAE model to ensure robustness and applicability across various asset classes.
@@ 3 Methodology
In this study, we employ a two-stage model to predict quantiles of asset returns. The first stage uses only asset-specific data to predict quantiles for normalized returns. The second stage incorporates market data to generate a scaling factor, which is then multiplied by the output from the first stage. This scaling factor adjusts the normalized returns based on broader market conditions. To train the model, we utilize a modified quantile loss function that accommodates multiple observations, in line with the generalized approach. This allows for a more robust evaluation of the model’s predictive capability across different datasets. We validate the methodology using both real-world and synthetic data.
The code used to implement all approaches in this section as well as to evaluate the models is publicly available at https://github.com/izzak98/dist-pred
@@ 3.1 Problem Definition
$X=\begin{bmatrix}x_{1,1}&x_{1,2}&\cdots&x_{1,t}\\ x_{2,1}&x_{2,2}&\cdots&x_{2,t}\\ \vdots&\vdots&\ddots&\vdots\\ x_{m,1}&x_{m,2}&\cdots&x_{m,t}\\ \end{bmatrix}$
Accompanying the asset matrix is the market data for the same $t$ time steps, represented as a matrix $Z\in\mathbb{R}^{g\times t}$, where $g$ is the number of market features, that is
@@
Our goal is to predict the distribution of log-returns of the asset for the future time steps from $t+1$ to $t+n$, where $n$ represents both the number of time steps used for the predictions and the forecast horizon. The vector of log returns $\vec{r}\in\mathbb{R}^{n}$ is defined as:
$\vec{r}=\begin{bmatrix}r_{t+1}\\ r_{t+2}\\ \vdots\\ r_{t+n}\\ \end{bmatrix}$
where each $r_{t+k}$ represents the log return at time step $t+k$ for $k=1,\dots,n$.
Given the input features $X$ and $Z$, the objective is to model and predict the distribution of future log returns
@@
where $f(X,Z)$ represents the predictive model for the log returns based on the asset and market features.
@@ 3.2 Model Architecture
To model the distribution of future log returns, we aim to predict the quantiles $\vec{\tau}$ of the return distribution over the prediction horizon. Utilizing kernel density estimation these can later be converted into distributions. The quantile vector is defined as
$\vec{\tau}=\begin{bmatrix}\tau_{1}\\ \tau_{2}\\ \vdots\\ \tau_{\mathcal{T}}\\ \end{bmatrix}$
where $\mathcal{T}$ denotes the number of quantiles, the quantiles used can be found in F.
For this task, we utilize Long Short-Term Memory (LSTM) networks, a type of recurrent neural network (RNN) [16] that are well-suited for sequential data modeling due to their ability to capture long-term dependencies and are less prone to issues such as vanishing gradients. Unlike many other time-series models, LSTMs can flexibly handle variable-length sequences, making them appropriate for our task as we want a variable amount of sequence lengths.
Following the methodology outlined in [2], we adopt a two-stage approach to predict the distribution of future log returns.
3.2.1 Stage 1: Asset-Specific Quantile Prediction
In the first stage, the model focuses on predicting the quantiles of future normalized returns $\vec{\tilde{r}}$ by leveraging the asset-specific features $X$ and the average volatility $\bar{\sigma}_{t}$ of all assets within the same asset group. This is done by predicting the quantiles $Q_{\vec{\tilde{r}}}^{\tau}$ of the volatility-normalized returns $\vec{\tilde{r}}$, where
@@ 3.2.1 Stage 1: Asset-Specific Quantile Prediction
$\vec{r}=\begin{bmatrix}r_{t+1}\\ \vdots\\ r_{t+n}\end{bmatrix}\quad\textrm{and}\quad\vec{\bar{\sigma}}=\begin{bmatrix}\bar{\sigma}_{t+1}\\ \vdots\\ \bar{\sigma}_{t+n}\end{bmatrix}$
Here, $\vec{\bar{\sigma}}$ represents the average volatility of assets in the same group at time step $t+k$, for $k=1,\dots,n$, where $n$ is the prediction horizon.
The average volatility $\bar{\sigma}_{t}$ is calculated as the exponentially weighted moving average (EWMA) of the log returns across all $K$ assets in the group, with a decay factor of 0.94, in line with standard practices and the methodology from [2], formally:
@@
with $\lambda$ being the decay factor, in our case 0.94, and $\bar{\sigma}_{t}$ is calculated as:
@@
where $\sigma_{i,t}$ is the volatility of asset $i$ at time $t$. We can formally represent this stage as:
$f_{1}(X)=Q_{\vec{\tilde{r}}}^{\tau}$
@@ 3.2.2 Stage 2: Market Data Scaling
In the second stage, the model incorporates the broader market data features $Z$ to refine the predictions from the first stage. The market data is used to create a scaling factor $\vec{\tilde{\sigma}}$, which adjusts the asset-specific quantile predictions based on market conditions.
The scaling factor $\vec{\tilde{\sigma}}$ is generated by applying a transformation to the market feature matrix $Z$, capturing the influence of market conditions on future returns. Additionally, the average volatility $\bar{\sigma}_{t}$ of all assets within the same asset group is used to further scale the predicted quantiles. This results in the final quantile prediction for the asset’s log returns:
$Q_{\vec{r}}^{\tau}=\vec{\hat{\sigma}}\cdot\vec{\bar{\sigma}}\cdot Q_{\vec{\tilde{r}}}^{\tau}$
@@
$f_{2}(Z)=\vec{\hat{\sigma}}$
Giving the final model the representation:
$\vec{r}\sim f(X,Z)=f_{\text{KDE}}(f_{1}(X)\cdot\vec{\bar{\sigma}}\cdot f_{2}(Z)$
where $f_{\text{KDE}}$ represents the kernel density estimations, for more details see Appendix D.

### v2

# Generalized Distribution Estimation for Asset Returns

@@ Abstract
Accurately modeling the distribution of asset returns enables better risk management, portfolio optimization, and financial decision-making especially in uncertain and volatile markets. This paper presents a novel approach to estimating the distribution of asset log returns using quantile regression combined with smoothed density estimation. Our proposed model, qLSTM, leverages asset-neutral features to deliver robust and generalizable predictions across diverse asset classes, outperforming baseline and dense models in quantile loss metrics for real-world data. A hybrid model, averaging Gaussian and model-predicted quantiles, qHybrid, further improves performance on synthetic datasets with skewed and heavy-tailed distributions. Evaluation using metrics like Wasserstein distance, CRPS, and VaR demonstrates the models’ effectiveness in capturing tail risks and broader distributional characteristics.
@@
Distribution Estimation , Qauntile Regression , LSTM , Tail Risk , Financial Modeling , Machine Learning
journal: Applied Soft Computing
[…]
@@ 1 Introduction
In the financial world, accurate modeling of asset price movements and their associated risks is a cornerstone of decision-making for investors, portfolio managers, and risk analysts [1]. This modeling forms the foundation of derivative pricing, risk management, and portfolio optimization. One widely used approach assumes that asset price changes over time can be modeled as a stochastic process, with the distribution of returns playing a central role in these models [2].
Machine learning has transformed several practices and processes withing the finance domain by enabling models that capture the complex, nonlinear relationships inherent in market data [3]. Applications range from price and return prediction to risk management and portfolio optimization, leveraging methods such as ARIMA, LSTM, and CNNs [4, 5, 6]. However, much of this research focuses on point predictions rather than estimating the full return distribution, which is essential for understanding uncertainty and tail risks.
@@
Traditionally, logarithmic asset returns are often assumed to follow a normal distribution; a famous example is the Black-Scholes model, which is used to find the fair value of option derivatives [2]. However, in practice, this assumption frequently does not hold [7]. Returns often display non-normal characteristics, such as skewness or heavy tails. Consequently, models that rely on normality assumptions can lead to suboptimal or even incorrect results [8]. To address this issue, some approaches generalize assumptions or adopt alternative distributions, such as the Student’s $t$- distribution, which accommodates skewness, kurtosis, and heavy tails [9]. Furthermore, distributional approaches in machine learning, show recent advances in conditional modeling and quantile regression, address this gap by providing a stronger method for distribution estimation [10, 11].
More recently, a new approach has emerged, namely to model the distribution directly [12]. This approach offers several key advantages. A more accurate model of the return distribution could improve derivative pricing, inform risk management, and enhance portfolio optimization. Although the proposed method shows promising results, it only considers one day’s worth of data and has a fixed look ahead period of 22 days. Furthermore, the method relies heavily on derived features which are exclusive to stocks, meaning the method does not generalize to other asset classes.
@@
In this paper, we propose a model that estimates the distribution of returns given current observations. Our model predicts the quantiles of the log return distribution over $n$ days given $n$ days worth of data, thus extending previous work[12]. Although quantile regression is not a novel concept, we convert these quantiles into a full distribution using a smoothed density estimation function. This enables a more precise estimation of the return distribution over a given time period. The model is trained on a dataset of asset returns and general market variables, ensuring that each characteristic is asset-class neutral, which makes the approach applicable across different asset classes.
This paper contributes to the literature in the following ways:
  • 1.
    Proposes a generalizable distribution estimation model that:
    • (a)
      Uses only asset-neutral features.
    • (b)
      Allows the consideration of multiple periods of return.
  • 2.
    Proposes a hybrid model that utilizes standard Gaussian quantiles in combination with the generalized model.
  • 3.
    Introduces an experimental framework for enhanced reproducibility in the field by using only data which is is publicly available.
The results are compared in two ways: quantile loss is used to evaluate the performance of our approaches against a baseline linear quantile regression model and the original dense approach in [12]. We find that our proposed qLSTM model achieves the lowest overall quantile loss for real-world data, demonstrating its robustness and versatility across diverse asset classes, including cryptocurrencies, commodities, and equity indices. The hybrid model, qHybrid, performs particularly well on synthetic data, especially for distributions with heavy tails or skewness, highlighting its flexibility. Additionally, we evaluate the models using standard metrics such as the Wasserstein distance and the Continuous Ranked Probability Score (CRPS). These metrics reveal that qLSTM excels in capturing extreme tail risks, while the qHybrid model achieves better calibration for broader distributions.
The rest of this paper is structured as follows. In Section 2, we provide an overview of the existing literature on machine learning in finance, focusing on the state of machine learning in finance, distributional modeling, and quantile regression methods. Section 3 outlines our methodology, including the problem formulation, model architecture, and training procedure. The data sources, including real-world and synthetic datasets, are described in detail, along with feature engineering and hyperparameter optimization. Section 4 presents the results of our experiments, comparing the performance of our proposed models with existing baselines using various metrics. Finally, Section 5 concludes the paper by summarizing our contributions and findings.
@@ 2 Literature Review
The application of machine learning to estimate phenomena within finance is well-explored, with much of the literature focusing on optimizing returns by predicting directions, prices, or returns. As noted by Dessain [13], the exponential growth in academic studies has made it increasingly difficult to synthesize existing knowledge within the domain. Similarly, Strader et al. [14] and Shah et al.. [15] emphasize the diversity of methods, such as ARIMA [4, 16], LSTM [5, 17], CNN [6, 18], and hybrid models [19, 20], as well as the inconsistency in the benchmarks, which complicates the comparisons between studies. To address these issues, our study uses publicly available data to ensure reproducibility and comparability.
In the domain of distribution estimation, significant advances have been made in non-conditional density estimation [21, 10, 22]. In contrast, this work focuses on conditional distributions allowing us to take various features from both the asset in question as well as features from the market to inform the shape of the distribution. Hu and Nan propose a method for estimating conditional distributions by modeling the hazard function and deriving the cumulative distribution function (CDF) via integration [11], while Feindt [23] demonstrates the utility of CDF-based methods across domains, including finance. However, both approaches prioritize global density estimation. Our work differs by directly estimating the quantile function, emphasizing financial applications, and incorporating domain-specific knowledge to capture tail risks.
Another notable contribution is by Rasul et al.. [10], who use normalizing flows conditioned on time series data for multistep distributional estimation, where the conditional distribution of $r_{k}$ depends not only on past observations $X$, but also on intermediate estimations $r_{1},\ldots,r_{k-1}$. This allows uncertainty to propagate across horizons. In contrast, we assume that each return on the forecast trajectory is drawn from the same distribution (i.e., $r_{1}\sim f(X),r_{2}\sim f(X)$), simplifying the computation while focusing on quantile estimation and tail risk for financial applications.
Focusing specifically on finance, Hickling and Prangle [24] leverage extreme value theory inspired transformations within normalizing flows to model heavy-tailed distributions. Similarly, Barun and Liu. [25] utilize quantile regression to address distributional challenges in financial returns. Building on this foundation, our work incorporates long-short-term memory (LSTM) networks to estimate quantiles over $n$ day horizons, ensuring robustness across asset classes by utilizing asset-neutral features, in contrast to feature-heavy approaches like those in [12], which rely on stock-specific characteristics and macroeconomic variables.
Despite progress, challenges remain in balancing model complexity, interpretability, and applicability across asset classes. This work addresses gaps in tail risk modeling by combining quantile-based LSTM architectures with asset-neutral features, providing robust and generalizable estimations across diverse financial datasets.
@@ 3 Methodology
Our proposed methodology consists of several steps. First we formalize the problem which this paper tackles, then we describe our two stage model to take into account both asset specific features as well as market general features. It is then followed by details about the two stage quantile loss function deployed. Next, we describe both real and synthetic data sources used in the training and evaluation of the models. Last, we detail the experimental setup and the metrics used for evaluation.
The code used to implement all the approaches in this section, as well as to evaluate the models, is publicly available at https://github.com/izzak98/dist-pred
@@ 3.1 Problem Definition
$X=\begin{bmatrix}x_{1,1}&x_{1,2}&x_{1,3}&\cdots&x_{1,t}\\ x_{2,1}&x_{2,2}&x_{2,3}&\cdots&x_{2,t}\\ x_{3,1}&x_{3,2}&x_{3,3}&\cdots&x_{3,t}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ x_{m,1}&x_{m,2}&x_{m,3}&\cdots&x_{m,t}\\ \end{bmatrix}$
where $x_{i,j}$ is the value of feature $i$ at time $j$. Accompanying the asset matrix is the market data for the same $t$ time steps, represented as a matrix $Z\in\mathbb{R}^{g\times t}$, where $g$ is the number of market features, that is
@@
where $z_{i,j}$ is value of market asset $i$ at time $j$.
Our goal is to estimate the distribution of the log returns of the asset for the future time steps from $1$ to $t$, where $t$ represents both the number of time steps used for the estimation and the forecast horizon. The vector of log returns $\vec{r}\in\mathbb{R}^{t}$ is defined as:
$\vec{r}=\begin{bmatrix}r_{1}\\ r_{2}\\ \vdots\\ r_{n}\\ \end{bmatrix}.$
Given the input features $X$ and $Z$, the objective is to model and estimate the distribution of log returns, that is,
@@
where $f(X,Z)$ represents the model for the estimation of log returns based on the asset and market features.
@@ 3.2 Model Architecture
To model the distribution of log returns, our objective is to estimate the quantiles $\vec{\tau}$ of the log return distribution over the estimation horizon. Utilizing density estimation, these can later be converted into distributions. The quantile vector is defined as
$\vec{\tau}=\begin{bmatrix}\tau_{1}\\ \tau_{2}\\ \vdots\\ \tau_{\mathcal{T}}\\ \end{bmatrix}$ (1)
where $\mathcal{T}$ denotes the number of quantiles. For the purpose of this research, we chose quantiles that represent broad ranging characteristics of a distribution, including quartiles and deciles as well as more fine grained intervals at the lower and upper end of the distribution to capture heavy tails. The quantiles used in this research can be found in A.
For this task, we use long- and short-term memory (LSTM) networks, a type of recurrent neural network (RNN) [26], which are well suited for sequential data modeling due to their ability to capture long-term dependencies and are less prone to issues such as vanishing gradients. LSTMs maintain two states at each time step: the cell state $\vec{c}_{t}$ and the hidden state $\vec{h}_{t}$, which are updated using gating mechanisms. In addition there are the input and output gates, $\vec{i}_{t}$ and $\vec{o}_{t}$, respectively. Specifically, the updates are governed by the equations:
$\vec{i}_{t}$ $=\sigma(W_{i}\vec{x}_{t}+U_{i}h_{t-1}+\vec{b}_{i}),$
$\vec{f}_{t}$ $=\sigma(W_{f}\vec{x}_{t}+U_{f}\vec{h}_{t-1}+\vec{b}_{f}),$
$\vec{o}_{t}$ $=\sigma(W_{o}\vec{x}_{t}+U_{o}\vec{h}_{t-1}+\vec{b}_{o}),$
$\tilde{\vec{c}}_{t}$ $=\tanh(W_{c}\vec{x}_{t}+U_{c}\vec{h}_{t-1}+\vec{b}_{c}),$
$\vec{c}_{t}$ $=\vec{f}_{t}\odot\vec{c}_{t-1}+\vec{i}_{t}\odot\tilde{\vec{c}}_{t},$
$\vec{h}_{t}$ $=\vec{o}_{t}\odot\tanh(\vec{c}_{t}),$
where $\sigma(\cdot)$ and $\tanh(\cdot)$ are the activation functions of the sigmoid and hyperbolic tangent, respectively, and $\odot$ denotes the multiplication of elements. Unlike many other time-series models, LSTMs can flexibly handle variable-length sequences, making them appropriate for our task as we are working with a variable number of sequence lengths.
Loosely following the methodology outlined in [12], we adopt a two-stage approach to estimate the distribution of future log returns.
3.2.1 Stage 1: Asset-Specific Quantile Estimation
In the first stage, the model focuses on estimating the quantiles of future normalized returns $\vec{\tilde{r}}$ by leveraging the asset-specific features $X$. This is done by estimating the quantiles $Q_{\vec{\tilde{r}}}^{\tau}$ of the volatility-normalized returns $\vec{\tilde{r}}$, where
@@ 3.2.1 Stage 1: Asset-Specific Quantile Estimation
$\vec{r}=\begin{bmatrix}r_{1}\\ \vdots\\ r_{t}\end{bmatrix}\quad\textrm{and}\quad\vec{\bar{\sigma}}=\begin{bmatrix}\bar{\sigma}_{1}\\ \vdots\\ \bar{\sigma}_{t}\end{bmatrix}$
Here, $\vec{\bar{\sigma}}$ represents the average volatility of assets in the same group at some time step, where $t$ is the estimation horizon.
The average volatility $\bar{\sigma}_{t}$ is calculated as the exponentially weighted moving average (EWMA) of the log returns across all $K$ assets in the group, with a decay factor of 0.94, in line with standard practices and the methodology from [12]. Formally the expression is
@@
where $\lambda=0.94$ is the decay factor. Here, $\bar{\sigma}_{t}$ is calculated as
@@
where $\sigma_{i,t}$ is the volatility of asset $i$ at time $t$. We represent this first stage with
$f_{1}(X)=Q_{\vec{\tilde{r}}}^{\tau},$
that is, a function of the input features gives the quantiles of the volatility-normalized log returns.
@@ 3.2.2 Stage 2: Market Data Scaling
In the second stage, the model incorporates the broader market data features $Z$ to refine the estimation from the first stage. The market data is used to create a scaling factor $\vec{\tilde{\sigma}}$, which adjusts the asset-specific quantile estimation based on market conditions.
The scaling factor $\vec{\hat{\sigma}}$ is generated by applying a transformation to the market feature matrix $Z$, capturing the influence of market conditions on future returns. This results in the final quantile estimation for the asset’s log returns:
$Q_{\vec{r}}^{\tau}=\vec{\hat{\sigma}}\cdot Q_{\vec{\tilde{r}}}^{\tau}$ (2)
where $Q_{\vec{r}}^{\tau}$ represents the final estimated quantile for the asset’s log returns, $\vec{\bar{\sigma}}$ is the scaling factor derived from the market data features $Z$, which adjusts the asset-specific estimations based on the broader market environment, and $Q_{\vec{\tilde{r}}}^{\tau}$ is the asset-specific quantile estimation from Stage 1, normalized by the asset’s volatility. We can formally represent this stage as
$f_{2}(Z)=\vec{\hat{\sigma}},$
that is, a function of the market data gives a scalar value to scale the quantiles from $f_{1}(X)$.
Now let $f_{\text{DE}}$ be a Quantile-Smooth Density Estimation. This method takes quantiles and their associated probabilities to generate a smooth probability density function (PDF) while ensuring that the CDF is monotonic, bounded, and normalized. It smooths the density to prevent spikes, enforces non-negativity, and normalizes the PDF to integrate to one, providing a well-behaved representation of the underlying distribution; for more details see the appendix LABEL:app:DE. This leads to the final model which is given by the representation
$\vec{r}\sim f(X,Z)=f_{\text{DE}}\left[f_{1}(X)\cdot f_{2}(Z)\right]$
where $f_{\text{DE}}$ represents the Quantile-Smooth Density Estimation. We refer to this model as qLSTM.
Hybrid model
We also propose a hybrid model, which we refer to as qHybrid, where after training the qLSTM model, we compute the average of the quantile of the model and a Gaussian fit of the log returns in the lookback period, $-t$ to 00 used in $X$. Formally, we define this as
$Q_{\mathcal{H}}^{\tau}=\frac{Q_{\mathcal{N}}^{\tau}+Q_{\vec{\tilde{r}}}^{\tau}}{2},$ (3)
where $Q_{\mathcal{N}}^{\tau}$ represents the Gaussian quantiles across the $\vec{\tau}$ defined in Eq. (1), computed as
$Q_{\mathcal{N}}^{\tau}=[\mu+\sigma\Phi^{-1}(\tau_{i})]_{i=1}^{\mathcal{T}},$
with $\Phi(x)$ being the standard normal cumulative distribution function:
@@
Here, $\mu$ and $\sigma$ denote the mean and standard deviation of the observed returns in the lookback period, from $-t$ to 00.